### 17 Apr 2026 v1

@@ 4.2 Goodness-of-fit testing
We draw repeated observations of the right-hand side of (4.45) based on the eigenvalues from the estimator $W^{1/2}\widehat{\Gamma}W^{1/2}$ to simulate the distribution of $\lVert G_{n,T}\rVert_{L_{2}}^{2}$ under $\mathcal{H}_{0}$, from which an appropriate quantile can be extracted. [^12] Since $\widehat{\Gamma}$ is an estimator of a covariance matrix, it can possess negative eigenvalues. We follow Andersen et al. (2024) and retain only those terms in the sum associated with positive eigenvalues.
@@ 5 Simulation analysis
In the “continuous-time” version of the model, we set $\Delta=1/N$ with N=23,400N=23{,}400. We assume data is captured over a time interval $[0,T]$ with $T=50,100,\dots,2000$, but the log-price process is only observed at a much coarser equidistant grid $t_{i}=i/n$ with $n=39,78,390$, for $i=0,1,\dots,nT$. This design is intended to imitate a security that is traded in a 6.5 hour window each day and whose price is recorded at the ten-, five-, and one-minute frequency over at most an eight-year period. This resembles the dataset analyzed in our empirical application. To estimate the latent spot volatility, associated with each $n$ we employ a bandwidth of $h_{n}=36,48,120$. With the above interpretation, this amounts to a six-, four-, and two-hour window. Hence, as $n$ increases we collect a larger amount of high-frequency data while reducing the time span of the estimation window, as stipulated by the rate conditions in the asymptotic theory. These choices follow Li et al. (2013) and Christensen et al. (2019). [^15] In unreported results that are available at request, we conducted a more extensive set of numerical experiments with other bandwidth choices. This showed that the average squared estimation error of the spot volatility estimator was basically flat over an extended region, so long as $h_{n}$ was not too small or too large.
@@ 6 Empirical application
Note. Ticker is the short name of the futures contract, $N$ is average daily number of transactions (in 1000s), whereas Min., 1Q, Median, Mean, 3Q, Max., Skewness, and Kurtosis is for the distribution of the log-realized variance over the whole sample.
@@
In what follows, we study the log-spot realized variance, which is less influenced by extreme outliers. We remark that by the invariance principle the copula of a random vector does not change under a strictly increasing transformation, so the realized copula of volatility is unaffected by studying the log-transform. In Figure 4, we show a relative frequency histogram for the log-realized variance of ES in Panel A, whereas Panel B is for TY. In agreement with the sample average and quantiles reported in Table 2, we observe that volatility in the stock market is, on average, higher than in the bond market. Furthermore, and consistent with Andersen et al. (2001, 2003), the shapes are close to the Gaussian bell curve, although both exhibit a mild positive skewness and excess kurtosis compared to the normal distribution. As evident from the right-hand side of Table 2, this effect is more pronounced in TY than ES.
To provide an initial assessment on the degree of volatility codependency in these time series, Figure 5 reports the average daily log-spot realized variance for ES in Panel A and TY in Panel B. The smoothing attenuates the measurement error in the spot volatility estimator a bit to better capture any comovement. There is forceful evidence of a positive relationship, most notably during periods of elevated market distress. To support this claim, in Panel A of Figure 6 we draw a scatter plot of the log-realized variance of ES (on the $x$-axis) against the log-realized variance of TY (on the $y$-axis). The graph reinforces the previous impression, which is further backed by Pearson’s and Kendall’s sample correlation coefficient. [^21] Kendall’s tau has been transformed as $\hat{\rho}_{\text{K}}^{a}=\sin(\pi/2\hat{\rho}_{\text{K}}^{u})$, where $a$ ($u$) denotes the adjusted (unadjusted = raw) estimator. The correction—motivated by Greiner’s equality—makes it unbiased for the correlation coefficient of a random sample from a bivariate normal distribution and, hence, its scale is comparable to the Pearson’s measure.
@@
In one end of the domain, $L$ is convergent toward $\lim_{z\rightarrow 1^{-}}L(z)=1$, whereas $R$ has the limit $\lim_{z\rightarrow 0^{+}}R(z)=1$. This holds for every copula, and so it is not very informative. In the other end, however, $L(z)$ and $R(z)$ can converge to any number in the interval $[0,1]$. Provided the limits exist, the lower and upper tail concentration parameter are defined as $\lambda_{L}=\lim_{z\rightarrow 0^{+}}L(z)$ and $\lambda_{U}=\lim_{z\rightarrow 1^{-}}R(z)$. If either limit is nonzero, we say there is asymptotic dependence in that tail. Furthermore, noting that $L(0.5)=R(0.5)$, it is common to plot $T(z)=\min(L(z),R(z))$ as the aggregate tail concentration function.
@@
To assess goodness-of-fit, we examine a composite null hypothesis with the Gumbel, Clayton, and Independence copula as contenders for describing the data. The Clayton and Independence copula are both heavily rejected with a $P\text{-value}$ close to zero, however. Conversely, the Gumbel copula is more appropriate with a $P\text{-value}=0.1136$.
@@ 7 Conclusion
We start by proposing a multivariate extension of the realized distribution function of volatility by Li et al. (2013), from which the realized copula of volatility is derived. Under in-fill asymptotics with a fixed time span, we then show that it affords a consistent estimator of the empirical copula of the latent stochastic volatility (measuring the codependency of volatility on the part of the sample path observed so far). In a double-asymptotic framework with the long-span dimension also going to infinity, our estimator converges to the stationary marginal copula of volatility, provided it exists. We derive a functional central limit theorem for the realized-copula process, which supports uniform inference on the volatility copula. To render the limit theory operational, we design a feasible estimator of the asymptotic covariance function. Based on this, we develop pointwise confidence intervals, uniform confidence bands, and a goodness-of-fit test to evaluate a hypothesis about the marginal copula of volatility.
@@
The asymptotic theory developed here for the latent volatility of semimartingale processes and the associated inference based on local realized proxies of spot variance can be extended in several directions in future research. First, the baseline analysis can be pursued in the presence of market microstructure noise by building on the univariate treatment of the realized distribution function of volatility in Christensen et al. (2019). A second avenue is to allow the realized copula of volatility to vary with the state of the economy or with observable macro-financial indicators, so that dependence in volatility can differ across regimes or along the business cycle. A third extension is to embed the volatility copula in a higher-dimensional framework together with a model for time-varying correlations, thereby delivering a unified copula-based representation of the dynamics of the full covariance matrix. These developments require additional structure but can be built directly on the econometrics and inference procedures developed in this paper.
@@ Appendix A Appendix
The following lemma, which extends Lemma 1 of Li et al. (2013) to the multivariate setting, gives the consistency for the estimation of a general functional of the variance processes.
@@
We denote $k_{n}\Delta_{n}=u_{n}$ and set $\hat{V}_{s}^{Z+}=\hat{V}_{\lceil{\frac{s}{u_{n}}}\rceil u_{n}}^{Z+}$, for $Z=X,Y$. By Lemma 1 of Li et al. (2013), it follows that $\hat{V}_{s}^{Z+}\overset{\mathbb{P}}{\longrightarrow}V_{s}$ for all $s\in\big(0,(\lfloor\frac{T}{u_{n}}\rfloor-1)u_{n}\big)$, delivering that $(\hat{V}_{s}^{X+},\hat{V}_{s}^{Y+})\overset{\mathbb{P}}{\longrightarrow}(V_{s}^{X},V_{s}^{Y})$. Thus, by condition (ii) and the Bounded Convergence Theorem (BCT), for Lebesgue almost everywhere $s\in[0,T]$, we get
@@
This is a bivariate version of Theorem 1 of Li et al. (2013) and Theorem 3.1 of Christensen et al. (2019).
@@
Therefore, we get $\mathbb{P}(V_{t}^{X}=x,V_{t}^{Y}=y)=0$ for Lebesgue almost everywhere for $t\in[0,T]$, which establishes condition (ii) in Lemma 1. The pointwise consistency of $\widehat{H}_{n,T}(x,y)$, namely $\widehat{H}_{n,T}(x,y)\overset{\mathbb{P}}{\longrightarrow}H_{T}(x,y)$, follows from Theorem 1 of Li et al. (2013) and Lemma 1. Since $\widehat{H}_{n,T}$ and $H_{T}$ are bounded and non-decreasing, pointwise convergence also implies locally uniform convergence.
@@
Part (2): By Corollary 1 in Li et al. (2013), which builds on Lemma 21.2 in van der Vaart (1998), we deduce that
@@
The proof of Theorem 2 relies on the upcoming lemma, which shows that the estimation error of the realized distribution function of volatility and the realized copula of volatility, centered around their empirical counterpart and scaled with $\sqrt{T}$, converges to zero in probability as $\Delta_{n}\rightarrow 0$ and $T\rightarrow\infty$. It extends Lemma 2 and Theorem 2 of Li et al. (2013) for $T$ fixed.
@@
Following Lemma 2 and Theorem 2 of Li et al. (2013), we know that $\eta_{n}^{X}$ and $\beta_{n}^{X}$ are both of order $O_{P}(d_{n})$, with
@@
To handle the second term in the inequality, namely $\sup_{t\in[0,\infty)}|F_{T}(\hat{V}_{t}^{X})-F_{T}(V_{t}^{X})|$, we follow the proof of Lemma 2(b) in Li et al. (2013) by observing that
@@
Reusing Li et al. (2013, Lemma 5), we deduce that
@@
Consider the mapping $\phi(H)(u,v)=H(Q^{X}(u),Q^{Y}(v))$. By Fermanian et al. (2004, Lemma 2), the mapping $\phi$ is Hadamard differentiable at the point $H^{\ast}(\cdot,\cdot\cdot)=H(Q^{X}(\cdot),Q^{Y}(\cdot\cdot))$. Moreover, the Hadamard derivative of $\phi$ at this point is given by

### 21 Apr 2026 v2

@@ 4.2 Goodness-of-fit testing
We draw repeated observations of the right-hand side of (4.45) based on the eigenvalues from the estimator $W^{1/2}\widehat{\Gamma}W^{1/2}$ to simulate the distribution of $\lVert G_{n,T}\rVert_{L_{2}}^{2}$ under $\mathcal{H}_{0}$, from which an appropriate quantile can be extracted. [^12] Since $\widehat{\Gamma}$ is an estimator of a covariance matrix, it can possess negative eigenvalues. We follow Andersen, Su, Todorov, and Zhang (2024) and retain only those terms in the sum associated with positive eigenvalues.
@@ 5 Simulation analysis
In the “continuous-time” version of the model, we set $\Delta=1/N$ with N=23,400N=23{,}400. We assume data is captured over a time interval $[0,T]$ with $T=50,100,\dots,2000$, but the log-price process is only observed at a much coarser equidistant grid $t_{i}=i/n$ with $n=39,78,390$, for $i=0,1,\dots,nT$. This design is intended to imitate a security that is traded in a 6.5 hour window each day and whose price is recorded at the ten-, five-, and one-minute frequency over at most an eight-year period. This resembles the dataset analyzed in our empirical application. To estimate the latent spot volatility, associated with each $n$ we employ a bandwidth of $h_{n}=36,48,120$. With the above interpretation, this amounts to a six-, four-, and two-hour window. Hence, as $n$ increases we collect a larger amount of high-frequency data while reducing the time span of the estimation window, as stipulated by the rate conditions in the asymptotic theory. These choices follow Li, Todorov, and Tauchen (2013) and Christensen, Thyrsgaard, and Veliyev (2019). [^15] In unreported results that are available at request, we conducted a more extensive set of numerical experiments with other bandwidth choices. This showed that the average squared estimation error of the spot volatility estimator was basically flat over an extended region, so long as $h_{n}$ was not too small or too large.
@@ 6 Empirical application
Note. Ticker is the short name of the futures contract, $N$ is the average daily number of transactions (in 1000s), whereas Min., 1Q, Median, Mean, 3Q, Max., Skewness, and Kurtosis are for the distribution of the log-realized variance over the whole sample.
@@
In what follows, we study the log-spot realized variance, which is less influenced by extreme outliers. We remark that by the invariance principle the copula of a random vector does not change under a strictly increasing transformation, so the realized copula of volatility is unaffected by studying the log-transform. In Figure 4, we show a relative frequency histogram for the log-realized variance of ES in Panel A, whereas Panel B is for TY. In agreement with the sample average and quantiles reported in Table 2, we observe that volatility in the stock market is, on average, higher than in the bond market. Furthermore, and consistent with Andersen, Bollerslev, Diebold, and Ebens (2001); Andersen, Bollerslev, Diebold, and Labys (2003), the shapes are close to the Gaussian bell curve, although both exhibit a mild positive skewness and excess kurtosis compared to the normal distribution. As evident from the right-hand side of Table 2, this effect is more pronounced in TY than ES.
To provide an initial assessment on the degree of volatility codependency in these time series, Figure 5 reports the average daily log-spot realized variance for ES in Panel A and TY in Panel B. The smoothing attenuates the measurement error in the spot volatility estimator a bit to better capture any comovement. There is forceful evidence of a positive relationship, most notably during periods of elevated market distress. To support this claim, in Panel A of Figure 6 we draw a scatter plot of the log-realized variance of ES (on the $x$-axis) against the log-realized variance of TY (on the $y$-axis). The graph reinforces the previous impression, which is further backed by Pearson’s and Kendall’s sample correlation coefficients. [^21] Kendall’s tau has been transformed as $\hat{\rho}_{\text{K}}^{a}=\sin(\pi/2\hat{\rho}_{\text{K}}^{u})$, where $a$ ($u$) denotes the adjusted (unadjusted = raw) estimator. The correction—motivated by Greiner’s equality—makes it unbiased for the correlation coefficient of a random sample from a bivariate normal distribution and, hence, its scale is comparable to the Pearson’s measure.
@@
At one end of the domain, $L$ is convergent toward $\lim_{z\rightarrow 1^{-}}L(z)=1$, whereas $R$ has the limit $\lim_{z\rightarrow 0^{+}}R(z)=1$. This holds for every copula, and so it is not very informative. In the other end, however, $L(z)$ and $R(z)$ can converge to any number in the interval $[0,1]$. Provided the limits exist, the lower and upper tail concentration parameters are defined as $\lambda_{L}=\lim_{z\rightarrow 0^{+}}L(z)$ and $\lambda_{U}=\lim_{z\rightarrow 1^{-}}R(z)$. If either limit is nonzero, we say there is asymptotic dependence in that tail. Furthermore, noting that $L(0.5)=R(0.5)$, it is common to plot $T(z)=\min(L(z),R(z))$ as the aggregate tail concentration function.
@@
To assess goodness-of-fit, we examine a composite null hypothesis with the Gumbel, Clayton, and Independence copulas as contenders for describing the data. The Clayton and Independence copulas are both heavily rejected with a $P\text{-value}$ close to zero, however. Conversely, the Gumbel copula is more appropriate with a $P\text{-value}=0.1136$.
@@ 7 Conclusion
We start by proposing a multivariate extension of the realized distribution function of volatility by Li, Todorov, and Tauchen (2013), from which the realized copula of volatility is derived. Under in-fill asymptotics with a fixed time span, we then show that it affords a consistent estimator of the empirical copula of the latent stochastic volatility (measuring the codependency of volatility on the part of the sample path observed so far). In a double-asymptotic framework with the long-span dimension also going to infinity, our estimator converges to the stationary marginal copula of volatility, provided it exists. We derive a functional central limit theorem for the realized-copula process, which supports uniform inference on the volatility copula. To render the limit theory operational, we design a feasible estimator of the asymptotic covariance function. Based on this, we develop pointwise confidence intervals, uniform confidence bands, and a goodness-of-fit test to evaluate a hypothesis about the marginal copula of volatility.
@@
The asymptotic theory developed here for the latent volatility of semimartingale processes and the associated inference based on local realized proxies of spot variance can be extended in several directions in future research. First, the baseline analysis can be pursued in the presence of market microstructure noise by building on the univariate treatment of the realized distribution function of volatility in Christensen, Thyrsgaard, and Veliyev (2019). A second avenue is to allow the realized copula of volatility to vary with the state of the economy or with observable macro-financial indicators, so that dependence in volatility can differ across regimes or along the business cycle. A third extension is to embed the volatility copula in a higher-dimensional framework together with a model for time-varying correlations, thereby delivering a unified copula-based representation of the dynamics of the full covariance matrix. These developments require additional structure but can be built directly on the econometrics and inference procedures developed in this paper.
@@ Appendix A Appendix
The following lemma, which extends Lemma 1 of Li, Todorov, and Tauchen (2013) to the multivariate setting, gives the consistency for the estimation of a general functional of the variance processes.
@@
We denote $k_{n}\Delta_{n}=u_{n}$ and set $\hat{V}_{s}^{Z+}=\hat{V}_{\lceil{\frac{s}{u_{n}}}\rceil u_{n}}^{Z+}$, for $Z=X,Y$. By Lemma 1 of Li, Todorov, and Tauchen (2013), it follows that $\hat{V}_{s}^{Z+}\overset{\mathbb{P}}{\longrightarrow}V_{s}$ for all $s\in\big(0,(\lfloor\frac{T}{u_{n}}\rfloor-1)u_{n}\big)$, delivering that $(\hat{V}_{s}^{X+},\hat{V}_{s}^{Y+})\overset{\mathbb{P}}{\longrightarrow}(V_{s}^{X},V_{s}^{Y})$. Thus, by condition (ii) and the Bounded Convergence Theorem (BCT), for Lebesgue almost everywhere $s\in[0,T]$, we get
@@
This is a bivariate version of Theorem 1 of Li, Todorov, and Tauchen (2013) and Theorem 3.1 of Christensen, Thyrsgaard, and Veliyev (2019).
@@
Therefore, we get $\mathbb{P}(V_{t}^{X}=x,V_{t}^{Y}=y)=0$ for Lebesgue almost everywhere for $t\in[0,T]$, which establishes condition (ii) in Lemma 1. The pointwise consistency of $\widehat{H}_{n,T}(x,y)$, namely $\widehat{H}_{n,T}(x,y)\overset{\mathbb{P}}{\longrightarrow}H_{T}(x,y)$, follows from Theorem 1 of Li, Todorov, and Tauchen (2013) and Lemma 1. Since $\widehat{H}_{n,T}$ and $H_{T}$ are bounded and non-decreasing, pointwise convergence also implies locally uniform convergence.
@@
Part (2): By Corollary 1 in Li, Todorov, and Tauchen (2013), which builds on Lemma 21.2 in van der Vaart (1998), we deduce that
@@
The proof of Theorem 2 relies on the upcoming lemma, which shows that the estimation error of the realized distribution function of volatility and the realized copula of volatility, centered around their empirical counterpart and scaled with $\sqrt{T}$, converges to zero in probability as $\Delta_{n}\rightarrow 0$ and $T\rightarrow\infty$. It extends Lemma 2 and Theorem 2 of Li, Todorov, and Tauchen (2013) for $T$ fixed.
@@
Following Lemma 2 and Theorem 2 of Li, Todorov, and Tauchen (2013), we know that $\eta_{n}^{X}$ and $\beta_{n}^{X}$ are both of order $O_{P}(d_{n})$, with
@@
To handle the second term in the inequality, namely $\sup_{t\in[0,\infty)}|F_{T}(\hat{V}_{t}^{X})-F_{T}(V_{t}^{X})|$, we follow the proof of Lemma 2(b) in Li, Todorov, and Tauchen (2013) by observing that
@@
Reusing Li, Todorov, and Tauchen (2013, Lemma 5), we deduce that
@@
Consider the mapping $\phi(H)(u,v)=H(Q^{X}(u),Q^{Y}(v))$. By Fermanian, Radulović, and Wegkamp (2004, Lemma 2), the mapping $\phi$ is Hadamard differentiable at the point $H^{\ast}(\cdot,\cdot\cdot)=H(Q^{X}(\cdot),Q^{Y}(\cdot\cdot))$. Moreover, the Hadamard derivative of $\phi$ at this point is given by